# SLAM3R: Real-Time Dense Scene Reconstruction from Monocular RGB Videos

### v2

@@ 4 Experiments
Table 1: Reconstruction results on 7 Scenes [49] dataset. We report accuracy and completeness in centimeters. The best results are shown in bold.The average numbers are computed over all test sequences.
@@
NICER-SLAM [77] 2.53 / 3.04 3.93 / 4.10 3.40 / 3.42 5.49 / 6.09 3.45 / 4.42 4.02 / 4.29 3.34 / 4.03 3.03 / 3.87 3.65 / 4.16 $\ll$1
DROID-SLAM [54] 12.18 / 8.96 8.35 / 6.07 3.26 / 16.01 3.01 / 16.19 2.39 / 16.20 5.66 / 15.56 4.49 / 9.73 4.65 / 9.63 5.50 / 12.29 $\sim$20
@@
Table 2: Reconstruction results on Replica [52] dataset. We report accuracy and completeness in centimeters. The best results are shown in bold. * denotes the results reported in NICER-SLAM.
@@
Datasets.
For both the Image-to-Points (I2P) and Local-to-World (L2W) models, we perform training with a mixture of three datasets: ScanNet++ [69], Aria Synthetic Environments [3] and CO3D-v2 [39]. These datasets vary from scene-level to object-centric, and contain both real-world and synthetic scenes. Since they are all recorded sequentially, we can easily extract video clips with a sliding-window mechanism as our training data. We select about 880K clips for training in total. To validate our reconstruction quality, we conduct quantitative evaluations on two unseen datasets: 7 Scenes [49], a real-world dataset of partial scenes, and Replica [52], a synthetic dataset of complete scenes. We also demonstrate visual reconstruction results across diverse datasets and in-the-wild captured videos to showcase the generalization ability of SLAM3R.
Implementation details.
Both the I2P and L2W models build upon the architecture of DUSt3R [62] with minimal but effective modifications, making it natural for them to initialize their weights from the DUSt3R pre-trained model. We initialize our weights using the DUSt3R model trained on 224$\times$224 resolution images with $m=24$ encoder blocks, $n=12$ decoder blocks with linear heads.
To train our I2P model, we extend the training process of DUSt3R from two views to multiple views. Specifically, our I2P model takes as input a video clip of length 11 with a suitable stride between adjacent frames, and designates the middle frame as the keyframe. We train the I2P model for 100 epochs, which takes about 6 hours.
After that, we train the retrieval module built on the I2P model. During training, we freeze all other modules and use L1 loss to supervise the correlation score against the mean confidence of the I2P model’s final predictions. This module requires only 20 epochs of training.
The window size is set to 13 when training the L2W model. A confidence-aware loss without scale normalization is applied, ensuring that the predicted point maps retain consistent scale with the input scene frames. The model is trained for 200 epochs, with the entire training process taking approximately 15 hours.
Due to computational constraints, all training images are center-cropped to 224 × 224 pixels. Standard data augmentation techniques [62] are applied. Our training is conducted on 8 NVIDIA 4090D GPUs, each with 24 GB of memory. For more details please refer to our supplementary material.
@@ 4.1 Comparisons
Following the implementation of NICER-SLAM [77] and Spann3R [59], we build a ground-truth point cloud model for each test scene by back-projecting pixels to the world using ground-truth depths and camera parameters. To address scene scale issues, we initially align our dense pointmap predictions with the ground truth using Umeyama algorithm [56]. This alignment is further refined through an ICP [41] procedure, ensuring optimal correspondence between the predicted and ground-truth point clouds. Reconstruction quality is quantified by computing the accuracy and completeness of the two point clouds. Furthermore, we report the frames per second (FPS) achieved during the reconstruction process on a single NVIDIA 4090D GPU to demonstrate the computational efficiency of our approach.
Experiments on the 7 Scenes [49] dataset.
@@
The numerical results of scene reconstruction quality are reported in Table 1. Following Spann3R [59]’s setting, we uniformly sample one-twentieth of the frames in each test sequence as input video. Each video is regarded as an individual scene. We evaluate SLAM3R using two settings: directly using the full pointmaps predicted for all input frames to create reconstruction results (denoted by SLAM3R-NoConf), and filtering each frame’s predicted pointmaps with a simultaneously predicted confidence map before creating reconstruction results (SLAM3R). We compare our method with optimization-based reconstruction DUSt3R [62], triangulation-based MASt3R [26], and online incremental reconstruction Spann3R. DUSt3R is tested using the weight-224 model with a resolution of 224$\times$224, the same as our input resolution, while MASt3R is tested using the weight-512 model with a resolution of 512$\times$384. Our method outperforms all baselines in both accuracy and completeness while maintaining real-time performance. As shown in the Office-09 scene (the top row in Figure 4), our approach demonstrates much less drift compared to the concurrent work Spann3R.
Experiments on the Replica [52] dataset.
@@
Besides the baselines mentioned in 7 Scene datasets, we also compare the SLAM-based reconstruction approaches NICER-SLAM [77], DROID-SLAM [54], DIM-SLAM [27] and GO-SLAM [73] on the Replica [52] dataset. The numerical results on full scene reconstruction are reported in Table 2. Due to the memory constraint, DUSt3R [62] and MASt3R [26] process only one-twentieth of the frames for reconstruction. As is shown in the table, our method surpasses all baselines with FPS greater than 1. Notably, without any optimization procedure, our method achieves reconstruction quality comparable to optimization-based methods such as NICER-SLAM [77] and DUSt3R [62]. Example of the Office 2 (the bottom row in Figure 4) also illustrates the global consistency of our reconstruction result.
@@ 4.2 Analyses
Effectiveness of the I2P model.
To highlight the advantages of our multi-view I2P model over the original two-view DUSt3R [62], we evaluate the reconstruction quality of keyframes with varying numbers of supporting views. We conduct experiments on the Replica [52] dataset, where input views are sampled using a sliding window of different sizes, and the reconstruction accuracy and completeness of the keyframes are computed. The results are reported in Table 3. As the number of supporting views increases, our approach progressively improves reconstruction quality. Notably, the efficiency of our method remains stable until the window size exceeds 11, demonstrating the effectiveness of our parallel design. However, the results also show diminishing returns as the number of views increases, which we detail in the supplementary material. Visual results of I2P reconstruction can be found in Figure 1.
@@
Table 3: Inner-window keyframe reconstruction results with various window lengths.
@@
Advantages of the L2W model.
The effectiveness of the L2W model is evaluated through ablation studies on the Replica [52] dataset. Per-window reconstructions are first generated with a window size of 11 using the I2P model. Local points are then aligned to a unified coordinate frame using different methods: global optimization from DUSt3R [62] (I2P-GA), traditional approaches such as Umeyama [56] and ICP [41] (I2P+UI), and our L2W model (I2P+L2W+Re). For consistency, we set the window size for global optimization to 10, which is equal to the number of views used to align new frames in other methods. Results in Table 4 show that our full method achieves superior alignment accuracy and computational efficiency compared to the alternatives.
Analysis of the retrieval module.
we propose a lightweight retrieval module that selects historical scene frames from the reservoir. This approach effectively performs implicit re-localization. We compare our retrieval method with a baseline approach that selects the ten nearest previous frames, named I2P+L2W. The results in Table 4 indicate a significant performance improvement with our retrieval strategy, demonstrating its effectiveness.
In-the-wild scene reconstruction.
We have tested our method on a diverse range of unseen datasets and found that SLAM3R shows strong generalization capabilities. Figure 5 shows our reconstruction results on Tanks and Temples [24], BlendedMVS [67], Map-free Reloc [2], LLFF [33], and ETH3D [48, 47] datasets, as well as in-the-wild videos we captured. These results show that our method performs reliably across different scales and scenes.
@@ 5 Conclusion
In this paper, we present SLAM3R, a novel and efficient monocular RGB SLAM system for real-time high-quality dense 3D reconstruction. It employs a two-hierarchy neural network framework to perform end-to-end 3D reconstruction through streamlined feed-forward processes, eliminating the need to explicitly solve any camera parameters. Experiments demonstrate its state-of-the-art reconstruction quality and real-time efficiency, achieving 20+ FPS.
6 Limitation and Future Work
Our system reconstructs scenes without explicitly solving or optimizing camera parameters. While this approach is efficient, it prevents us from performing global bundle adjustment. Consequently, in large-scale outdoor scenes, we still face the challenge of accumulated drift. Addressing this limitation will be a focus of our future work.
@@ Appendix A Implementation details
$\mathcal{L}_{Retr}$ $=\sum_{i=1}^{R}\lvert S_{i}^{\prime}-Mean(C_{i}^{\prime})\rvert$
$S_{i}^{\prime}$ $=Sigmoid(S_{i})$
$C_{i}^{\prime}$ $=(C_{i}-1)/C_{i}$
@@
To construct the training data, we utilize all iPhone and DSLR frames registered by COLMAP [45] from the training and validation splits of ScanNet++[69]. Additionally, we include all frames from the first 450 scenes of the Aria Synthetic Environments (ASE)[3] dataset and 41 categories from CO3D-v2 [39], with each category containing up to 50 randomly sampled scene sequences. We introduce two ways to extract video clips for training. For ScanNet++ and ASE, we adopt uniform sampling with strides of 3 and 2, respectively. For CO3D-v2, frames are randomly sampled within temporal segments covering half the length of each video. In total, we extract approximately 880K clips. During each epoch of training for the I2P model, we randomly sample 2000, 1000, and 1000 clips of length 11 from the ScanNet++, ASE, and CO3D-v2 datasets, respectively. When training the L2W model, the number of clips sampled per epoch doubles, and the clip length increases to 13. When training with ground truth pointmaps, we set invalid points to (0,0,0). The effective batch sizes of training are 16 for the I2P model and 32 for the L2W model.
@@
DUSt3R [62] 9.03 8.46 3.16 15.83 13.85 13.00 25.31 12.65 $<$1
DUSt3R [62] (gt intrinsic) 5.08 5.08 2.47 13.19 10.96 10.71 13.27 8.68 $<$1
MASt3R [26] 4.31 2.55 1.78 11.37 11.31 7.84 5.27 6.35 $\ll$1
MASt3R [26] (gt intrinsic) 4.68 2.91 1.43 13.94 12.82 8.21 3.18 6.74 $\ll$1
NICER-SLAM [77] 3.28 6.85 4.16 10.84 20.00 3.94 10.81 8.55 $<$1
DIM-SLAM [27]* 60.13 2.41 39.42 24.15 13.07 39.11 10.45 26.96 $\sim$3
DROID-SLAM [54] 3.36 2.40 1.43 9.19 16.46 4.94 1.85 5.66 $\sim$20
DROID-SLAM [54]* 3.55 2.49 3.32 10.93 48.53 4.72 2.58 10.87 $\sim$20
Spann3R [59] 9.86 6.83 7.02 23.85 12.87 15.16 13.89 12.78 $>$50
SLAM3R (Ours) 6.59 5.99 12.24 15.52 15.05 11.44 12.19 11.27 $\sim$25
Table 5: Camera pose estimation results on the 7Scenes [49] dataset reported using the ATE-RMSE (cm) metric. The average numbers are computed over all test scenes.
@@
DUSt3R [62] (1/20) 7.69 8.28 11.41 7.00 4.82 5.16 5.32 8.19 7.23 $<$1
DUSt3R [62] (gt intrinsic) 3.97 4.80 7.65 5.10 3.99 3.84 2.95 6.42 4.84 $<$1
MASt3R [26] 1.60 1.88 4.05 1.57 6.72 1.49 75.09 5.09 12.19 $\ll$1
MASt3R [26] (gt intrinsic) 1.06 1.05 0.85 0.95 7.52 1.47 2.01 1.84 2.09 $\ll$1
NICER-SLAM [77] 1.36 1.60 1.14 2.12 3.23 2.12 1.42 2.01 1.88 $<$1
@@
DIM-SLAM [27]* 1.06 0.49 0.32 0.43 0.26 0.65 0.55 3.69 0.93 $\sim$3
DROID-SLAM [54] 0.34 0.13 0.27 0.25 0.42 0.32 0.52 0.40 0.33 $\sim$20
DROID-SLAM [54]* 0.58 0.58 0.38 1.06 0.40 0.70 0.53 1.33 0.70 $\sim$20
Spann3R [59] (1/20) 30.14 39.39 32.48 37.65 27.06 45.02 57.11 42.68 38.94 $>$50
SLAM3R (Ours) 4.62 5.95 5.46 10.96 7.18 5.85 5.27 16.33 7.70 $\sim$24

### v3

@@ 4 Experiments
Table 1: Reconstruction results on 7 Scenes [51] dataset. The average numbers are computed over all test sequences. The methods are categorized into two groups based on whether their FPS is above or below 1. The best results within each category are shown in bold. We report accuracy and completeness in centimeters. The color gradient shifts from red through yellow to green to show increasing FPS.
@@
NICER-SLAM [79]* 2.53 / 3.04 3.93 / 4.10 3.40 / 3.42 5.49 / 6.09 3.45 / 4.42 4.02 / 4.29 3.34 / 4.03 3.03 / 3.87 3.65 / 4.16 $\ll$1
DROID-SLAM [56]* 12.18 / 8.96 8.35 / 6.07 3.26 / 16.01 3.01 / 16.19 2.39 / 16.20 5.66 / 15.56 4.49 / 9.73 4.65 / 9.63 5.50 / 12.29 $\sim$20
@@
Table 2: Reconstruction results on Replica [54] dataset. * denotes the results reported in NICER-SLAM.
DUSt3R [64] MASt3R [28] NICER-SLAM [79]* DROID-SLAM [56]* DIM-SLAM [29] GO-SLAM [75] Spann3R [61] SLAM3R-NoConf (Ours) SLAM3R (Ours)
7 Scenes 8.02 6.28 8.55 5.66 - - 11.70 8.44 8.41
Replica 4.76 1.67 1.88 0.33 0.46 0.39 32.79 6.61 6.61
Table 3: Camera pose results evaluated by ATE-RMSE (cm) on 7 Scenes [51] and Replica [54] datasets.
@@
Datasets. For both the Image-to-Points (I2P) and Local-to-World (L2W) models, we perform training with a mixture of three datasets: ScanNet++ [71], Aria Synthetic Environments [3] and CO3D-v2 [41]. These datasets vary from scene-level to object-centric, and contain both real-world and synthetic scenes. Since they are all recorded sequentially, we can easily extract video clips with a sliding-window mechanism as our training data. We select about 850K clips for training in total. To validate our reconstruction quality, we conduct quantitative evaluations on two unseen datasets: 7 Scenes [51], a real-world dataset of partial scenes, and Replica [54], a synthetic dataset of complete scenes. We also demonstrate visual reconstruction results across diverse datasets and in-the-wild captured videos to showcase the generalization ability of SLAM3R.
Implementation details. Both the I2P and L2W models build upon the architecture of DUSt3R [64] with minimal but effective modifications, making it natural for them to initialize their weights from the DUSt3R pre-trained model. We initialize our weights using the DUSt3R model trained on 224$\times$224 resolution images with $m=24$ encoder blocks, $n=12$ decoder blocks with linear heads. All images are center-cropped to 224$\times$224 pixels before feeding into SLAM3R. Our training is conducted on 8 NVIDIA 4090D GPUs, each with 24 GB of memory. It takes about one day. At test time, we set the initial window length to $L=5$ to ensure high-quality reconstruction of all frames within the window. For subsequent incremental windows, we use $L=11$ to provide more supporting views for better keyframe reconstruction. Please refer to our supplementary material for more implementation details.
@@ 4.1 Comparisons
Evaluation metrics. Following NICER-SLAM [79] and Spann3R [61], we build a ground-truth point cloud model for each test sequence by back-projecting pixels to the world using ground-truth depths and camera parameters. The reconstructed point clouds are aligned to ground truths using Umeyama [58] and ICP [43] algorithms. We quantify reconstruction quality through accuracy and completeness metrics. To demonstrate computational efficiency, we report frames per second (FPS) on a single NVIDIA 4090D GPU. We also evaluate camera poses using absolute trajectory error (ATE-RMSE). For detailed formulations of the metrics, please refer to the supplementary material.
@@
Reconstruction results on the 7 Scenes [51] dataset. The numerical results of scene reconstruction quality are reported in Table 1. Following Spann3R [61]’s setting, we uniformly sample one-twentieth of the frames in each test sequence as input video. Each video is regarded as an individual scene. We evaluate SLAM3R using two settings: integrating the full pointmaps predicted for all input frames to create reconstruction results (denoted by SLAM3R-NoConf), and filtering pointmaps with a confidence threshold of 3 before creating reconstruction results (SLAM3R). We compare our method with optimization-based reconstruction DUSt3R [64], triangulation-based MASt3R [28], and online incremental reconstruction Spann3R. Our method outperforms all baselines in both accuracy and completeness while maintaining real-time performance. As shown in the Office-09 scene (the top row in Figure 4), our approach demonstrates much less drift compared to the concurrent work Spann3R [61].
@@
Reconstruction results on the Replica [54] dataset. Besides the baselines mentioned in 7 Scene datasets, we also compare the SLAM-based reconstruction approaches NICER-SLAM [79], DROID-SLAM [56], DIM-SLAM [29] and GO-SLAM [75] on the Replica [54] dataset. The numerical results on full scene reconstruction are reported in Table 2. Due to the memory constraint, DUSt3R [64] and MASt3R [28] process only one-twentieth of the frames for reconstruction. As is shown in the table, our method surpasses all baselines with FPS greater than 1. Notably, without any optimization procedure, our method achieves reconstruction quality comparable to optimization-based methods such as NICER-SLAM [79] and DUSt3R [64]. Example of the Office 2 (the bottom row in Figure 4) also illustrates the global consistency of our reconstruction result.
Camera pose estimation on 7 Scenes [51] and Replica [54]. Our method is designed in a new paradigm that reconstructs 3D points end-to-end without explicitly solving camera parameters. Following DUSt3R [64], We also derive camera poses from the predicted scene points using PnP-RANSAC solver in OpenCV [6] with ground truth camera intrinsics of each frame. The results are reported in Table 3. We can observe that camera poses and scene reconstruction results are not fully positively correlated. This discrepancy between pose and reconstruction errors indicates that effective end-to-end 3D reconstruction is possible and promising without first obtaining precise camera poses.
For more details on the comparisons in this section, please refer to our supplementary material.
@@ 4.2 Analyses
Effectiveness of the I2P model. To highlight the advantages of our multi-view I2P model over the original two-view DUSt3R [64], we evaluate the reconstruction quality of keyframes with varying numbers of supporting views. We conduct experiments on the Replica [54] dataset, where input views are sampled using a sliding window of different sizes, and the reconstruction accuracy and completeness of the keyframes are computed. The results are reported in Table 4. As the number of supporting views increases, our approach progressively improves reconstruction quality. Notably, the efficiency of our method remains stable until the window size exceeds 11, demonstrating the effectiveness of our parallel design. However, the results also show diminishing returns as the number of views increases, which we detail in the supplementary material. Visual results of I2P reconstruction can be found in Figure 1.
@@
Table 4: Inner-window keyframe reconstruction results with various window lengths. By default, we use 11-frame windows for incremental reconstruction to balance quality and efficiency.
@@
Advantages of the L2W model. The effectiveness of the L2W model is evaluated through ablation studies on the Replica [54] dataset. Per-window reconstructions are first generated with a window size of 11 using the I2P model. Local points are then aligned to a unified coordinate frame using different methods: global optimization from DUSt3R [64] (I2P-GA), traditional approaches such as Umeyama [58] and ICP [43] (I2P+UI), and our L2W model (I2P+L2W+Re). For consistency, we set the window size for global optimization to 10, which is equal to the number of views used to align new frames in other methods. Results in Table 5 show that our full method achieves superior alignment accuracy and computational efficiency compared to the alternatives.
Analysis of the retrieval module. We propose a lightweight retrieval module that selects historical scene frames from the reservoir. This approach effectively performs implicit re-localization. We compare our retrieval method with a baseline approach that selects the ten nearest previous frames, named I2P+L2W. The results in Table 5 indicate a significant performance improvement with our retrieval strategy, demonstrating its effectiveness.
In-the-wild scene reconstruction. We have tested our method on a diverse range of unseen datasets and found that SLAM3R shows strong generalization capabilities. Figure 5 shows our reconstruction results on Tanks and Temples [26], BlendedMVS [69], Map-free Reloc [2], LLFF [35], and ETH3D [50, 49] datasets, as well as in-the-wild videos we captured. These results show that our method performs reliably across different scales and scenes. We also provide additional numerical results on sampled scenes from these datasets in the supplementary material.
@@ 5 Conclusion
In this paper, we present SLAM3R, a novel and effective system that performs real-time, high-quality, dense 3D scene reconstruction using RGB videos. It employs a two-hierarchy neural network framework to perform end-to-end 3D reconstruction through streamlined feed-forward processes, eliminating the need to explicitly solve any camera parameters. Experiments demonstrate its state-of-the-art reconstruction quality and real-time efficiency (20+ FPS).
Limitations and future work. The elimination of camera parameter prediction prevents us from performing global bundle adjustment. Additionally, the poses derived from our scene point cloud prediction still fall short of SLAM systems that specialize in camera localization. Addressing these limitations will be the focus of our future work.
6 Acknowledgments
This work is supported by the National Key R&D Program of China (2022ZD0160800). This work is also supported by the Early Career Scheme of the Research Grants Council (grant # 27207224), the HKU-100 Award, a donation from the Musketeers Foundation, and in part by the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We thank the reviewers for their valuable feedback, as well as Zihan Zhu and Songyou Peng for their help with our experiments. Siyan Dong would also like to thank the support from HKU School of Computing and Data Science.
@@ Appendix A Implementation details
$\mathcal{L}_{Retr}$ $=\sum_{i=1}^{R}\lvert S_{i}^{\prime}-Mean(C_{i}^{\prime})\rvert,$
$S_{i}^{\prime}$ $=Sigmoid(S_{i}),$
$C_{i}^{\prime}$ $=(C_{i}-1)/C_{i},$
@@
To construct the training data, we utilize all iPhone and DSLR frames registered by COLMAP [47] from the training splits of ScanNet++[71]. Additionally, we include all frames from the first 450 scenes of the Aria Synthetic Environments (ASE)[3] dataset and 41 categories from CO3D-v2 [41], with each category containing up to 50 randomly sampled scene sequences. We introduce two ways to extract video clips for training. For ScanNet++ and ASE, we adopt uniform sampling with strides of 3 and 2, respectively. For CO3D-v2, frames are randomly sampled within temporal segments covering half the length of each video. In total, we extract approximately 850K clips. During each epoch of training, we randomly sample 4000, 2000, and 2000 clips from the ScanNet++, ASE, and CO3D-v2 datasets, respectively. All training images are resized and then center-cropped to 224 × 224 pixels. Standard data augmentation techniques [64] are applied.
To train our I2P model, we extend the training process of DUSt3R from two views to multiple views. Specifically, our I2P model takes as input a video clip of length 11, and designates the middle frame as the keyframe. We train the I2P model for 100 epochs, which takes about 6 hours. After that, we train the retrieval module built on the I2P model. During training, we freeze all other modules and use L1 loss to supervise the correlation score against the mean confidence of the I2P model’s final predictions. This module requires 50 epochs of training, which takes about 2 hours.
To train the L2W model, we use clips of length 12, with the first six images selected as scene frames, and the last six images designated as keyframes to register. The model is trained for 200 epochs in total, and the training process takes approximately 16 hours. When training with ground truth pointmaps as input, we set invalid points to (0,0,0). A confidence-aware loss without scale normalization is applied, ensuring that the predicted point maps retain consistent scale with the input scene frames.
Our training is conducted on 8 NVIDIA 4090D GPUs, each with 24GB of memory and a batch size of 4 per GPU.
@@
DUSt3R [64] (w/PnP) 5.09 4.88 2.52 12.07 10.64 10.35 10.55 8.02 $<$1
MASt3R [28] (w/PnP) 4.32 2.92 1.47 12.37 11.82 7.98 3.04 6.28 $\ll$1
NICER-SLAM [79]* 3.28 6.85 4.16 10.84 20.00 3.94 10.81 8.55 $<$1
DROID-SLAM [56]* 3.36 2.40 1.43 9.19 16.46 4.94 1.85 5.66 $\sim$20
Spann3R [61] 9.18 6.69 7.10 21.56 12.83 14.06 10.43 11.70 $>$50
SLAM3R-NoConf (Ours) 6.29 5.33 4.47 12.42 11.74 9.53 9.30 8.44 $\sim$25
SLAM3R (Ours) 6.20 5.30 4.56 12.40 11.71 9.47 9.20 8.41 $\sim$25
Table 6: Camera pose estimation results on the 7Scenes [51] dataset reported using the ATE-RMSE (cm) metric. The average numbers are computed over all test scenes. * denotes the results reported in NICER-SLAM.
@@
DUSt3R [64] (w/PnP) 4.00 4.49 7.62 4.88 4.04 3.90 2.84 6.30 4.76 $<$1
MASt3R [28] (w/PnP) 1.07 0.99 0.87 0.90 4.90 1.21 1.77 1.63 1.67 $\ll$1
NICER-SLAM [79]* 1.36 1.60 1.14 2.12 3.23 2.12 1.42 2.01 1.88 $<$1
@@
DIM-SLAM [29] 0.48 0.78 0.35 0.67 0.37 0.36 0.33 0.36 0.46 $\sim$3
DROID-SLAM [56]* 0.34 0.13 0.27 0.25 0.42 0.32 0.52 0.40 0.33 $\sim$20
Spann3R [61] 29.76 34.78 26.08 34.50 22.65 34.47 42.24 37.84 32.79 $>$50
SLAM3R-NoConf (Ours) 4.54 5.89 5.73 11.17 6.32 6.15 4.99 8.05 6.61 $\sim$24
SLAM3R (Ours) 4.56 5.88 5.72 11.17 6.32 6.15 4.95 8.09 6.61 $\sim$24